OpenStudy (anonymous):

Power Series. Prove that sum (n/(n+1)!) = 1

OpenStudy (anonymous):

\[\sum_{0}^{\infty} (n/(n+1)!)=1\]

OpenStudy (anonymous):

\[\sum_{0}^{\infty}\frac{ n+1 }{ (n+1)! } -\sum_{0}^{\infty}\frac{ n }{ (n+1)! } = e - 1\]